OpenStudy (anonymous):

is the line through the points R(-1,3) and S (2,-7) parallel to the graph of the line given by the equation 10x+3y=6

OpenStudy (helder_edwin):

find the slope between R and S. if it is the same as the slope of the given line, then they r parallel

OpenStudy (anonymous):

a.no, the slopes have opposite signs.b.yes, the slopes have the same sign and value. c.yes, the lines bith decrease to the right. d.no, the lines have unequal slopes

OpenStudy (helder_edwin):

do u know how to compute the slope between two points?

OpenStudy (anonymous):

no, sorry

OpenStudy (helder_edwin):

ok given R(a,b) and S(c,d) then the slope is given by \[\large m=\frac{b-d}{a-c} \] try it.

OpenStudy (anonymous):

sorry, I just wouldn't know how

OpenStudy (helder_edwin):

ok u have R(-1,3) and S(2,-7) then the slope is \[\large m=\frac{3-(-7)}{-1-2}=\frac{3+7}{-1-2}=\frac{10}{-3}=-\frac{10}{3} \]

OpenStudy (helder_edwin):

got it?

OpenStudy (anonymous):

I will keep studying this, thank you :)

OpenStudy (anonymous):

so b then right?

OpenStudy (helder_edwin):

the next step would be to get y in terms of x. u have \[\large 10x+3y=6 \] from which \[\large y=-\frac{10}{3}x+2 \] and the slope of this line is -10/3

OpenStudy (helder_edwin):

u r right, it is B

OpenStudy (anonymous):

thank you :)

OpenStudy (helder_edwin):

u r welcome